A child fires a foam dart from a dart gun. The dart initially travels with a speed of 6.00 m/s horizontal to the ground. The dart strikes the ground at a horizontal distance of 2.00 m from where the dart was fired. How high above the ground was the dart when it was fired?
Answer #1

The dart undergoes a projectile motion with an initial velocity of 6m/s only in the x direction.

Ux=6 m/s

Therefore here, we use the 2nd equation of motion to-

1) Calculate the time of travel in the x direction

In the ne dénection ото SQM Um Gm/s 2 s= ut + late 2) t=8-2 2 sec

2) Calculate the height in y direction using the time of travel

In the yairection, apg=908 m/s2 uy=0 - n=lyt that =48908 x 512=0.54m CtxI=0.54m. 3 3